OpenStudy (anonymous):

Is the expression x^3•x^3•x^3 equivalent to x^3•3•3? Why or why not? Explain your reasoning

OpenStudy (anonymous):

No, 'cause x^3.x^3.x^3 = x^(3+3+3)= x^9 [ laws of exponents, a^b .a^c =a^(b+c)]